Anna’s demand for peaches is given by PA = 200 – 3QA, where QA is the quantity (kilograms) demanded at price PA($/kilo)...

  1. Anna’s demand for peaches is given by PA = 200 – 3QA, where QA is the quantity (kilograms) demanded at price PA($/kilo). Basil’s demand is given by PB = 120 – 2QB, where QB is the quantity (kilograms) demanded at price PB ($/kilo). Each of the two has $2,000 that they can spend if they want to buy something.
    1. Suppose the endowments are as follows: Anna has 110 kilograms of peaches, Basil has none. When they trade, who will sell to whom? How many kilos will be sold? At what price?
    2. Suppose the endowments are as follows: Anna has no peaches, Basil has 110 kilos. When they trade, who will sell to whom? How many kilos will be sold? At what price?
    3. Suppose the endowments are as follows: Anna has 60 kilograms of peaches, Basil has 50 kilos. When they trade, who will sell to whom? How many kilos will be sold? At what price?
Answer #1

demamd PA 200-39A Apmnas 200-PA demand Basil s P=120-20B R 60-PB 2 niome Ama (MA)= 2000 Income Rasl (MB) 2000 1s EndocomentPh PR P20 Loth mdnidualy Prce foced Demomd by each peiso QA 5 3 200 20 60 20 S0 1I0 endonment , She Snce eAnna ha and Ranl wotndooment of Anna o 20 Endowmeut of Basil110 TOel endowmnt 10 to tio d r To endosmend TO ta demomd &A taR= g Fom paut ca) oeTuis condihon is exactly same au it uay m paut a So, after. trede pce amd quanhty grver by pautCa)-e are Theke an eractly the